OpenStudy (anonymous):

Nicole Is Taking Some Friends On A Picnic. They'll Need To Follow A Path To Get To Picnic Spot. A Map Of The Path Is Based On A Scale Of 1:40,000 In CM. If The Path Is 15 CM On The Map, How Many Kilometers Is The Actual Path 15 x 40,000=600,000 1 KM=400,000 CM 600,000/400,000=1.5

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

1 km = 100000cm can you write the whole question

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

im here can you write the whole quesiton

OpenStudy (anonymous):

I Did

OpenStudy (anonymous):

15 x 40,000=600,000 ---- correct 1 KM=400,000 CM ---- wrong 1 km = 100000cm 600,000/400,000=1.5 ---- fix the the above and you will get it

OpenStudy (anonymous):

@DamnImBad did you get it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But It Said It Was Incorrect And I Was Wondering How Come Its A Extra Zero On The 1 KM=100,00(0)

OpenStudy (anonymous):

1 km = 1,000 m 1 m = 100 cm 1,000 * 100 = 100,000

OpenStudy (anonymous):

But It Told Me It Was Incorrect But Thanks For The Help

OpenStudy (anonymous):

15 x 40,000=600,000 1 KM=100,000 CM 600,000/100,000 = 6

OpenStudy (anonymous):

Oh Now I C Y I Get It Wrong Thanks

OpenStudy (anonymous):

just make sure you pay attention when converting km to cm